I need help with a story problem.
A diamond is purchased for $2500. Suppose it increases 5% in value each year. Find the value after 8 years.

Thanks.

Found 2 solutions by stanbon, dolly:
Answer by stanbon(75887)

A diamond is purchased for $2500. Suppose it increases 5% in value each year. Find the value after 8 years.
==============
Each year its value is multiplied by (1 + 0.05) = value + 5%(value)
2500(1.05)^8 = $3693.64
=========
Cheers,
Stan H.

Answer by dolly(163)

Value of the diamond = $2500
Increases by 5% each year.
Value after 8 years = 2500(1 + 0.05)^8
= 2500(1.05)^8
= 2500(1.48)
= 3700
The diamond is worth $3700 after 8 years